Six litres of 0.4M 3-(N-morpholino) propane sulfonic acid (MOPS) buffer of pH 7.8 are required. Available in the laboratory are 10M HCl and 10M NaOH and the zwitterionic form of MOPS (M.Wt = 210.3). The pKa for the MOPS zwitter ion is 7.2. Describe, showing the relevant calculations, the procedure for preparing the buffer . If 10ml of 10M HCl was added to 1 litre of the above buffer (question 9) what would be the resulting pH?
What would be the resulting pH if 10ml of 10M HCl were added to 1 litre of pure water?
1
Expert's answer
2020-09-21T06:33:28-0400

The following formula is used to calculate the pH of the given buffer

"pH=pK_a +lg(\dfrac{C_{BA}}{C_{HA}})"

"pK_a=7.2"

Rearranging the formula

"\dfrac{C_{BA}}{C_{HA}}=10^{{pH}-pK_a}\newline"

"=10^{7.8-7.2}=3.98"


The concentration of the buffer solution is "= 0.4M"

"C_{BA}+C_{HA}=0.4"

The concentration of MOPS and Na-salt required can be now evaluated

"3.98C_{HA}+C_{HA}=0.4\newline C_{HA}=\dfrac{0.4}{3.98+1}=0.08M"

"C_{BA}=0.4-0.08=0.32M"

Calculate the number of moles of HA and BA

"n(BA)=C_{BA}*V_{solution}\newline
0.32*6=1.92mol"


"n(HA)=C_{HA}*V_{solution}\newline 0.08*6=0.48mol"

The total number of moles of the solution is

"1.92+0.48=2.40mol"

Mass of MOPS="2.4*210.3=504.7g"

Volume of NaOH=1.92/10=0.192 L


Mix 504.7 g of MOPS with 0.192 L of 10M NaOH solution

"C_7H_{15}NO_4S\ +NaOH\ =C_7H_{14}NO_4SNa\ +H_2O"

In a flask, dissolve this mixture with distilled water up to 6L after the reaction is complete

"pH=Pk_A
+log(\dfrac{B}{A})"

"=7.2+log(0.32/0.08)\newline
=7.2+log4\newline=7.8"